# Interpolated Stochastic Interventions based on Propensity Scores, Target Policies and Treatment-Specific Costs

### v2

@@ Introduction
Table 1 presents the results for a sample size of $n=1000$ and averaged over $200$ repetitions. Results across all three setups demonstrate the superior performance of the proposed one-step estimators over the plug-in estimators. Under correct specification, the one-step estimator achieves near-zero bias and lower or comparable RMSE, particularly under policy $\pi^{\star}_{\delta}$. Under policy $\nu^{\star}_{\delta}$, lower bias is traded for a moderately higher RMSE when the target distribution $\nu$ is not concentrated but spread over the sample space instead. The robustness of the proposed one-step estimators is especially notable under model misspecification: when the outcome regression $Q$ is misspecified, the one-step estimator maintains low bias and RMSE in all setups, whereas the plug-in estimator exhibits increases in both metrics, most dramatically under $\nu^{\star}_{\delta}$. Similarly, with misspecification of the propensity score $\pi$, the one-step estimator substantially reduces bias and RMSE relative to the plug-in approach, though performance degrades more than with $Q$-misspecification. Across setups, $\widehat{\mu}^{T}_{\delta}$ tends to be more sensitive to misspecification than $\widehat{\mu}^{S}_{\delta}$, but the one-step correction consistently mitigates this sensitivity, highlighting its robustness and practical advantage in finite-sample, imperfect-model settings.
@@
Given that methylphenidate is the most widely prescribed stimulant, and that adverse effects such as sleep disturbances and weight loss are commonly reported (Graham and Coghill 2008), the overall cost of treatment may encompass both direct healthcare expenditures and costs associated with managing adverse effects. To account for this, we evaluate two cost/target configurations, both assigning a higher cost to high-dose treatment: (a) a target in which 75% of individuals receive low-dose treatment and 25% receive high-dose treatment, and (b) a balanced target with a 50–50% allocation between low and high doses.
@@ Conclusions
We propose a cost-aware family of stochastic interventions for discrete treatments that generalizes incremental propensity score interventions and connects causal modeling to cost-sensitive decisions. Modeling a cost-penalized I-projection (CPIP) of the independent product of organic and target distributions yields closed-form Boltzmann–Gibbs couplings whose marginals, via a single tilt $\delta$, interpolate from the propensity scores or from the target policy toward a product-of-experts limit when destination costs are strictly positive. We derive efficient influence functions under a nonparametric model and construct one-step estimators with uniform bands over $\delta$, improving stability and misspecification robustness relative to plug-in baselines.
@@ Acknowledgments
This work was supported by the Research Council of Norway via grants with project numbers 302899 (Effects of pharmacological treatment and special education on school performance in children with ADHD) and 332645 (Integreat – Norwegian centre for knowledge-driven machine learning). Data curation for the application case was conducted by Dr. Guido Biele at the Norwegian Institute of Public Health. The study was approved by the Regional Committees for Medical and Healthcare Research Ethics (REK approval no. 96604).

### v3

@@ Introduction
Table 1 reports results for $n=1000$, averaged over $200$ Monte Carlo repetitions. Across all three setups, the proposed one-step estimators consistently outperform the corresponding plug-in estimators. Under correct specification, the one-step estimators attain essentially almost-zero bias and achieve lower or comparable RMSE, with the gains most pronounced after the policy $\pi^{\star}_{\delta}$. Under $\nu^{\star}_{\delta}$, the reduction in bias can come at the cost of a modest increase in RMSE when the target distribution $\nu$ is diffuse rather than concentrated on a single action. The robustness advantage is clearest when the outcome regression $\widehat{Q}$ is misspecified: the one-step estimators retain low bias and RMSE throughout, whereas the plug-in estimators incur substantial deterioration in both measures, especially after $\nu^{\star}_{\delta}$. When the propensity score model $\widehat{\pi}$ is misspecified, consistency is no longer theoretically guaranteed, yet the one-step correction still delivers lower bias and RMSE than plug-in across settings. Overall, $\widehat{\mu}^{T}_{\delta}$ is more sensitive to misspecification than $\widehat{\mu}^{S}_{\delta}$, but the one-step correction systematically attenuates this sensitivity, underscoring its practical value in finite samples under imperfect nuisance modeling.
@@
Given that methylphenidate is the most widely prescribed stimulant, and that adverse effects such as sleep disturbances and weight loss are commonly reported (Graham and Coghill 2008), the overall cost of treatment may encompass both direct healthcare expenditures and costs associated with managing adverse effects. To account for this, we evaluate two cost and target configurations, both assigning a higher cost to high-dose treatment: (a) a target in which 75% of individuals receive low-dose treatment and 25% receive high-dose treatment, and (b) a balanced target with a 50–50% allocation between low and high doses.
@@ Conclusions
We propose a cost-aware family of stochastic interventions for discrete treatments that generalizes incremental propensity score interventions and connects causal modeling to cost-sensitive decisions. Modeling a cost-penalized I-projection of the independent product of organic law and reference / target policy yields closed-form Boltzmann–Gibbs couplings whose marginals, via a single tilt $\delta$, interpolate from either of the input distributions toward a product-of-experts limit when destination costs are strictly positive. We derive efficient influence functions under a nonparametric model and construct one-step estimators with uniform bands over $\delta$, improving stability and misspecification robustness relative to plug-in baselines.
@@ Acknowledgments
This work was supported by the Research Council of Norway via two grants with project numbers 302899 (Effects of pharmacological treatment and special education on school performance in children with ADHD) and 332645 (Integreat – Norwegian centre for knowledge-driven machine learning). Data curation for the application case was conducted by Dr. Guido Biele at the Norwegian Institute of Public Health. The study was approved by the Regional Committees for Medical and Healthcare Research Ethics (REK approval no. 96604).